OpenStudy (anonymous):

Solve. x2 = 10x – 20

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

Hey, baby. Probably the simplest way for you to do this problem is to simple move everything to one side of the equation to get x^2-10x+20 =0 and then use the quadratic formula.

OpenStudy (anonymous):

anyone want to do another?

OpenStudy (hba):

the other way is to break the middle term

OpenStudy (anonymous):

Solve. 3x2 = –15x + 21

OpenStudy (anonymous):

Hey, Ken. As awesome as it is that alpha will do this for you, let me show you how to solve it on your own.

OpenStudy (anonymous):

First, get everything to one side; 3x^2 +15x -21 = 0 Then we can use the quadratic formula. A=3 B=15 C=-21 Plug those into the formula \[\Large x = \frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (jiteshmeghwal9):

Take 10x-20 at L.H.S.as\[x^2-10x-20=0\]now it becomes quadratic Equation so\[-(-10) \pm \sqrt{(-10)^2-4(1)(-20)}\over 2(1)\]now solve it & gt ur answer:)

OpenStudy (jiteshmeghwal9):

\[10 \pm \sqrt{10+80}\over2\]\[10 \pm \sqrt{90}\over 2\]